OpenStudy (anonymous):

the numbers 7,24 and c are a pythagorean triple.what is the value of c A.31 B.27 C.24 D.25

OpenStudy (unklerhaukus):

a pythogorean is a set in integers \(a,b,c\) that solve \[a^2+b^2=c^2\] if \(a\) is 7 and \(b\) is 24, what is \(c\)?

OpenStudy (anonymous):

would i add or muiltplay

OpenStudy (unklerhaukus):

square each of 7, and 24 , then add them

OpenStudy (unklerhaukus):

finally take the square root of this to get c

OpenStudy (anonymous):

so 49+576=625

OpenStudy (unklerhaukus):

so what is c?

OpenStudy (anonymous):

625 ?

OpenStudy (unklerhaukus):

625 is c^2 take the square root of 625 to get c ,

OpenStudy (anonymous):

25 !!!

OpenStudy (unklerhaukus):

yeahs

OpenStudy (anonymous):

i have another question

OpenStudy (unklerhaukus):

what it is

OpenStudy (anonymous):

ok a square patio has an area of 50 square feet.to the nearest tenth of a foot what is the length of the patio

OpenStudy (unklerhaukus):

use The area of a square is equal to its side length squared as an equation ,[ with the side length as \(s\) and the area as \(A\)] \[A=s^2\] if you take the square root of both sides \[\sqrt A=s\]

OpenStudy (anonymous):

would u square 8

OpenStudy (unklerhaukus):

if the sides were 8 the are would be 8^2 = 64 which is too large

OpenStudy (unklerhaukus):

the area*

OpenStudy (unklerhaukus):

take the square root of fifty to find your side length

OpenStudy (unklerhaukus):

if the area was 49 , the sides would be √49 = 7. but this is a little bit too small